Toxic pollutants. The annual cost in dollars for removing
p% of the toxic chemicals from a town’s water supply is
given by the formula
C(p) = 500,00/ 100-p
.
a) Use the accompanying graph to estimate the cost for
removing 90% and 95% of the toxic chemicals.
b) Use the formula to find C(99.5) and C(99.9).
c) What happens to the cost as the percentage of
pollutants removed approaches 100%?
: Toxic pollutants. The annual cost in dollars for removing
p% of the toxic chemicals from a town’s water supply is
given by the formula
C(p) = 500,00/ 100-p
.
a) Use the accompanying graph to estimate the cost for
removing 90% and 95% of the toxic chemicals.
b) Use the formula to find C(99.5) and C(99.9).
c) What happens to the cost as the percentage of
pollutants removed approaches 100%?

Answer by checkley77(3848)

C(p) = 500,00[0]/(100-p) I'LL ASSUME YOU LEFT OUR THE LAST 0 IN 500,000 FIGURE.
C(90)=500,000/(100-90)
C(90)=500,000/10
C(90)=50,000 ANS.
C(95)=500,000/(100-95)
C(95)=500,000/5
C(95)=100,000 ANS.
C(99.5)=500,000/(100-99.5)
C(99.5)=500,000/.5
C(99.5)=1,000,000 ANS.
C(99.9)=500,000/(100-99.9)
C(99.9)=500,00/.1
C(99.9)=5,000,000 ANS.
AS P APPROACHES 100% THE COST APPROACHES INFINITY.